### v1

# Continuous and Reversible Electrical Tuning of Fluorescent Decay Rate via Fano Resonance

Abstract
Decay rate of an atomic or molecular dipole depends on the local density of optical states (LDOS) —Purcell effect—which can significantly be enhanced near a dielectric-core metal-shell nanoparticle (CSNP). On top of that, an auxiliary quantum object (QO) can introduce a Fano transparency in the plasmonic spectrum of the CSNP. Here, we show that an auxiliary QO, located at the hotspot of the CSNP, can modify the LDOS, hence the decay rate of an excited dipole. Moreover, by controlling the resonance of the auxiliary QO via an applied voltage, one can continuously and reversibly tune both the radiative and nonradiative decay rates of the dipole up to 2 orders-of-magnitude. This phenomenon emerges as an invaluable tool to implement in integrated quantum technologies, enabling realization of on-demand entanglement/single-photon sources, controlled execution of quantum gates and electrical-control of superradiant-like phase transitions. It also bears potential for application in super-resolution microscopy and surface-enhanced Raman spectroscopy (SERS).
I Introduction
Electromagnetic interaction between an atomic system and a closely located metal nanostructure (MNS) has provided a viable platform for researchers to investigate the nature of nanophotonic phenomena yielding series of notable innovations such as SERS, that allows for single molecule detection [1], single photon transistors realizing strong nonlinear interactions at the single-photon level [2], and cavity-free subwavelength confinement of optical fields using 0D-1D structures [3]. At the heart of the demonstrated innovations lies the Purcell effect [4] which defines the change in the lifetime of an atomic/molecular excited state. It takes place due to the modifications in the number of available optical states (local density of optical states, LDOS) emitter can couple.
This, in turn, results in the enhancement or suppression of the spontaneous emission rate of the emitter. The effect is prominent especially when the molecular system is located in a resonant cavity [5] or near a MNS [6] that modifies the electromagnetic environment. The modified decay rate is given by the change in the LDOS with respect to free space, such as =/ where is the free space decay rate. In the presence of an absorptive material, such as an MNS, enhanced molecular decay rate can take place both as photon emission and nonradiative photon transfer (absorption) into the MNS —apart from the intrinsic nonradiative losses within the emitter. The former demonstrates itself as fluorescence enhancement [7] while the latter is linked to optical losses.
Active control of Purcell effect is a key to enable switchable light emitting nanophotonic devices which represents a fast-growing field due to its direct contribution to telecommunication and integrated quantum technologies [8, 9, 10]. While there exist studies on the control of spontaneous emission via Purcell effect, they mostly rely on altering the cavity size/parameters representing a passive control scheme [11, 12, 13]. Cavity-dependent demonstrations of Purcell effect have attracted great attention due to the ability to tune the fluorescence emission spectra of molecules. However, it comes with certain limitations for active control of Purcell effect such as fixed cavity components and mirror spacings [14]. Although this limitation was partially solved within a tunable microcavity scheme, the applicability of the approach is still limited in the context of integrated circuits due to the requirement of immersion oils and CMOS incompatible elements [15].
Alternatively, some active methods to switch photoluminescence (PL) have also been demonstrated using various physical phenomena such as quantum confined stark effect[16, 17], electric field induced quenching [18], Förster resonant energy transfer (FRET) [19], and implementation of electroactive polymers in redox reaction [20, 21]. Although it was shown to be possible to switch fluorescence actively, these methods utilize quenching in order to turn off the fluorescence. That is, molecular excitation is almost completely transferred (lost) into the MNS and not kept in the emitter. Moreover, the modulation depths are poor and some of the studies report only a limited number of successive modulations due to hysteresis in the cyclic performance prohibiting a complete recovery between on and off states [19, 20, 21]. Furthermore, the response time of those applications reaches maximum millisecond regime which is not compatible with CPU clock speed range (GHz) and barely supports current display technology with frame rates (FPS) of 60-144 Hz. Despite the attempts to use reverse bias providing 200:1 contrast ratio and 300 ns response time [22], the usage of extremely high (3106 V/cm2) external field prohibits the technological use of reported methodologies.
Therefore, achieving continuous and reversible control within a compact device geometry that provides large modulation depths with (and over) CPU clock speed compatibility remains as a great challenge in active control of fluorescence for technology applications.
In this paper, we provide a proof-of-principle demonstration of active and reversible tuning of both radiative and nonradiative decay rates of a fluorescent molecule (FM). We electrically tune the LDOS by shifting the position of a Fano transparency (also known as Fano resonance) [23, 24, 25]. The resonance of the Fano transparency, induced by a quantum object (QO) residing at the CSNP hotspot (Fig. 1a), is tuned by shifting the level-spacing of the QO via an applied voltage. This way, the decay rates can be continuously tuned between =1–215 within a “picosecond” response time which is limited by the readout rate of the circuit operation (GB/s). We emphasize that the demonstrated system is fundamentally different from previously reported active PL switching and modulation effects that are based on screening of the PL for an external observer. By trapping the fluorescent state in the molecule (i.e., not quenching it), the nature of our switching process exhibits a crucial property with the ability to hold and release the interested fluorescent state by voltage application.
In our demonstrated system, the presence of a plasmonic SiO2/Ag core (SiO2) shell (Ag) nanoparticle (CSNP) boosts the LDOS of a FM (Fig. 1a). When an auxiliary QO, such as a quantum dot (QD) or a defect-center of resonance is positioned near the CSNP, it introduces a Fano transparency at [26]. Fano transparency is the plasmon-analog of electromagnetically-induced-transparency (EIT) [27]. Coupling of a dark mode or a QO to a bright plasmon mode introduces two paths for the absorption [28]. When the two paths work out of phase, absorption is canceled, and a transparency window appears in the spectrum of the MNS [29]. Therefore, an absorption and scattering dip appears at . In this work, by choosing about the transition frequency of the FM’s metastable (fluorescence) state (Fig. 1b), we show that the presence of the QO turns off the LDOS that used to be enhanced due to the presence of the CSNP (Fig. 2a). In other words, at turns off the extra LDOS occurring due to the presence of the CSNP. Additionally, it also turns off the nonradiative decay rate emerging due to the dissipation at the Ag shell (Fig. 2b).
More interestingly, as the QO’s resonance can be shifted [30, 31] via applied voltage (Fig. 3a), the decay rate patterns of Figs. 2a and 2b are also shifted electrically, enabling the continuous tuning of the decay rates (Figs. 3b and 3c). Thus, the LDOS, which is initially enhanced due to the CSNP, can also be tuned continuously via the applied voltage within sub-picosecond response time. We emphasize that a 20 meV resonance shift can tune the radiative decay rate between =1–215 continuously.
This method, being reversible and cavity-free, allows for nanofabrication approaches for the integration in quantum circuits and conventional Si-based readout circuits (see the Supplementary Material, SM). Compared to active electroluminescent counterparts, Fano resonance provides an incomparably large range of tuning corresponding to 215 times (21500%) modulation. It presents technologically compatible operation frequencies since the response time of the Fano resonance is of the order of picosecond and therefore the tuning rate is only limited by the speed of the solid-state device. We believe that the demonstrated method manifests itself as an invaluable tool that enables so-far unrealized implementations ranging from voltage-controlled superpositions of hyperfine states to on-demand single-photon imaging techniques.
II Results
Fig. 1a shows a sample configuration we have simulated as a proof-of-principle demonstration of the phenomenon. In our simulations we use a SiO2/Ag CSNP of diameter 105 nm where plasmonic losses limiting the use of MNSs in optoelectronics are compensated via optical gain incorporated in the core insulator (SiO2) region (100 nm in diameter) [32, 33]. The Ag-shell is a 5 nm thick layer that is deposited uniformly on the SiO2 core. An auxiliary QO of 40 nm diameter representing a densely implanted G-centers (or nitrogen-vacancy (NV) defect centers) [34, 35, 36] is located 5 nm to the Ag-shell and introduces the Fano transparency at . FM (dipole), polarized along the x-axis, is positioned on the opposite side at 15 nm distance. We calculated the LDOS for the radiative and nonradiative decay rates of the dipole (fluorescent molecule FM) using finite difference time domain method (FDTD) [37]. In the demonstrated configuration, we first show that Fano transparency can switch off the initially enhanced emission (radiative decay) rate of the FM. Fig. 2a shows the radiative decay spectrum of the dipole (FM) with and without the QO in the system. The QO residing at the hotspot of the CSNP turns off the CSPN induced enhancement in the LDOS. Decay rate drops from 215 to at . For other fluorescence wavelengths, decay rate suppression is weaker. One should note that decay of a molecule from its excited state always includes a quantum probabilistic nature [38]. The is to be understood as an expectation value.
In Fig. 2b, we observe a similar effect for the nonradiative decay rate. The presence of the QO also suppresses the absorption losses that the CSNP induces initially. The nonradiative decay is suppressed down to 5 from 65. We model the QO by a Lorentzian dielectric function [26, 28, 39] of resonance =766 nm, linewidth Hz and oscillator strength . We used the corresponding experimental dielectric functions for CSNP composed of Ag shell and core materials [40, 41]. To show the promises of our method, we further study the voltage-controlled continuous tuning of the decay rates. Application of an external electric field on the QO shifts the resonance (Fig. 3a).
Thus, the dashed curves in Figs. 2a and 2b, in particular the dips, also shift. The FDTD simulations show that the radiative and nonradiative decay rates can be tuned between =1–215 and =5–70, respectively, in our sample setup (Figs. 3b and 3c). Moreover, as the electrical shift in the resonance is reversible, modulation of and are also reversible. In Figs. 3b and 3c, the fluorescence is fixed at = 766 nm and is altered. Only a meV tuning of the QO resonance is sufficient for such a control noting that much larger shifts have already been reported even for 1V of applied potential [42].
Fig. 3a shows the schematic illustration of a possible device where the collection of the defect centers is modelled as the QO. Several experimental methods have been reported for the formation of G-centers in Si/ as well as NV centers in diamond, such as the incorporation of carbon rich Si crystals in reactive ion etching process [35], functionalization of Si surface with organic molecules [34] and hot ion implementation in diamond [36]. In our model, by representing a collection of G-centers as a QO, the external field is provided through the gate of a field-effect transistor (FET), where the active material that contains defect centers are biased through the source-drain channel. To provide the agreement between our model and the FET scheme, a step-like structure is employed merely to provide the spatial alignment of the defect centers to the CSNP. We note that an inhomogeneous broadening among the resonances of, e.g., G-centers results a similar broadening in the Fano transparency dip in Figs. 2a and 2b. In this case, even better control over relatively broader fluorescence spectra can be achieved.
Thus, we numerically demonstrate that we can actively control both the radiative and nonradiative decay rates of a FM for a narrow wavelength regime in a cavity-free environment. Our simulations were originally performed in vacuum therefore the dielectric environment of any possible device scheme (as shown in Figure 3a) would shift the plasmon resonances. This, however, won’t change the nature of the demonstrated phenomenon.
III Conclusion, Discussion and Outlook
We provide a proof-of-principle numerical demonstration for the voltage-tunability of decay rates. Our simulations —exact solutions of Maxwell equations— show 2-orders magnitude modulation in the decay rate. A QO turns off the CSNP induced LDOS at fluorescence = . When is shifted by an applied voltage (Fig. 3a), the schemes (Figs. 2a and 2b) for the decay rates also shift, enabling a continuous tuning of these rates in Figs. 3b and 3c.
The phenomenon addresses a wide range of applications as we are demonstrating a fundamental issue in light-matter interaction, that is electrical-tuning of LDOS. This provides an invaluable tool especially for integrated quantum technologies.The technique offers ultimate control over on-demand single-photon sources and entanglement generation in circuits. These also imply voltage-controlled quantum gate operations. The phenomenon of superradiant phase transition, which depends on the decay rate, now can be electrically turned on/off in an integrated circuit. Thus, normal and superradiant phase transitions between zero and a very strong entanglement state [43], can be employed to serve as a controllable collective entanglement resource in the integrated circuits. In addition, quantum batteries can become voltage-controlled based on the demonstrated phenomenon [44, 45].
Change in the LDOS not only alters the decay rates but can also tune coupling of a fixed intensity light to dipoles. This enables electrical tuning of light-matter interaction and stimulated transitions in ion-trap configurations. It is worth noting that the coefficients among two or more hyperfine states can be tuned very quickly without necessitating a change in the pump intensity of different polarizations [46]. For appropriate choice of MNS-shapes (i.e. helical MNSs with spin angular momentum), the phenomenon can be used to enhance/control LDOS selectively among different polarizations. Thus, the method provides a compact, voltage-tunable and fast preparation of single atom/ion states or collective states of their ensembles. For direct applications with the state-of-the-art ion traps, however, one also needs to take extra care of interaction of MNS with the trap potentials, though such potentials result only weak couplings with MNSs. The same features are also valid for the quantum gate operations.
In this work, we study proof-of-principle demonstration of the phenomenon in free space where the tuning starts at and ranges to 215. It is well studied that decay rates of FMs can be lowered down to values that are much smaller than , e.g., when they are in photonic crystal cavities [47]. Such setups, in principle, can provide electrical tuning of decay rates in terms of the multiples of much smaller base values . Demonstrated device structure can be achieved by various fabrication techniques and device construction schemes. (See the SM for details of reported fabrication steps.) Finally, the effect that we observe can be made more efficient by implementing more optimal setups.
We anticipate that the demonstrated technique can also be utilized in quantum optics for the on-demand generation of single-photons (or photon-pairs) and their synchronization with external systems such as a vibrating scanning probe microscopy (SPM) tip, or a confocal scanning laser microscope. Thanks to the ultimate control on the time gated activation of the emitter, the technique can also prove to be useful in performing spin-echo based quantum measurements/sensing applications such as quantum magnetometers [48] or quantum information processing [49]. Alternately, it can also serve as a voltage tunable band rejection filter for the radiative and nonradiative wavelengths falling into the corresponding Fano resonance linewidth. In general, the suitable execution of the demonstrated technique can form a base technology in the applications that require fine tuning of a specific fluorescence spectrum such as SERS and super-resolution microscopy.
Acknowledgements.
We gratefully thank Mehmet Emre Tasgin and Serkan Ates for discussions leading to this research. This research is conducted in the meetings that took place at Institute of Nuclear Sciences, Hacettepe University. E.O.P acknowledges the personal research fund (KAF) of Bilkent University. R.S. acknowledges support from TUBITAK No. 121F030 and 123F156.References
- Serafinelli et al. [2022] C. Serafinelli, A. Fantoni, E. C. Alegria, and M. Vieira, Plasmonic metal nanoparticles hybridized with 2D nanomaterials for SERS detection: A review, Biosensors 12, 225 (2022).
- Chang et al. [2007] D. E. Chang, A. S. Sørensen, E. A. Demler, and M. D. Lukin, A single-photon transistor using nanoscale surface plasmons, Nature Physics 3, 807 (2007).
- Akimov et al. [2007] A. Akimov, A. Mukherjee, C. Yu, D. Chang, A. Zibrov, P. Hemmer, H. Park, and M. Lukin, Generation of single optical plasmons in metallic nanowires coupled to quantum dots, Nature 450, 402 (2007).
- Purcell [1995] E. M. Purcell, Spontaneous emission probabilities at radio frequencies, in Confined Electrons and Photons: New Physics and Applications (Springer, 1995) pp. 839–839.
- Vahala [2003] K. J. Vahala, Optical microcavities, Nature 424, 839 (2003).
- Lakowicz [2005] J. R. Lakowicz, Radiative decay engineering 5: metal-enhanced fluorescence and plasmon emission, Analytical Biochemistry 337, 171 (2005).
- Camposeo et al. [2015] A. Camposeo, L. Persano, R. Manco, Y. Wang, P. Del Carro, C. Zhang, Z.-Y. Li, D. Pisignano, and Y. Xia, Metal-enhanced near-infrared fluorescence by micropatterned gold nanocages, ACS Nano 9, 10047 (2015).
- Lu et al. [2017] Y.-J. Lu, R. Sokhoyan, W.-H. Cheng, G. Kafaie Shirmanesh, A. R. Davoyan, R. A. Pala, K. Thyagarajan, and H. A. Atwater, Dynamically controlled purcell enhancement of visible spontaneous emission in a gated plasmonic heterostructure, Nature Communications 8, 1631 (2017).
- Munnix et al. [2009] M. C. Munnix, A. Lochmann, D. Bimberg, and V. A. Haisler, Modeling highly efficient rcled-type quantum-dot-based single photon emitters, IEEE Journal of Quantum Electronics 45, 1084 (2009).
- Bimberg et al. [2009] D. Bimberg, E. Stock, A. Lochmann, A. Schliwa, J. A. Tofflinger, W. Unrau, M. Munnix, S. Rodt, V. A. Haisler, A. I. Toropov, et al., Quantum dots for single-and entangled-photon emitters, IEEE Photonics Journal 1, 58 (2009).
- Canet-Ferrer et al. [2012] J. Canet-Ferrer, L. J. Martínez, I. Prieto, B. Alén, G. Muñoz-Matutano, D. Fuster, Y. González, M. L. Dotor, L. González, P. A. Postigo, et al., Purcell effect in photonic crystal microcavities embedding InAs/InP quantum wires, Optics Express 20, 7901 (2012).
- Bennett et al. [2005] A. Bennett, D. Unitt, P. See, A. Shields, P. Atkinson, K. Cooper, and D. Ritchie, Microcavity single-photon-emitting diode, Applied Physics Letters 86 (2005).
- Moritake et al. [2016] Y. Moritake, Y. Kanamori, and K. Hane, Emission wavelength tuning of fluorescence by fine structural control of optical metamaterials with fano resonance, Scientific Reports 6, 33208 (2016).
- Steiner et al. [2008] M. Steiner, A. V. Failla, A. Hartschuh, F. Schleifenbaum, C. Stupperich, and A. J. Meixner, Controlling molecular broadband-emission by optical confinement, New Journal of Physics 10, 123017 (2008).
- Chizhik et al. [2009] A. Chizhik, F. Schleifenbaum, R. Gutbrod, A. Chizhik, D. Khoptyar, A. J. Meixner, and J. Enderlein, Tuning the fluorescence emission spectra of a single molecule with a variable optical subwavelength metal microcavity, Physical Review Letters 102, 073002 (2009).
- Empedocles and Bawendi [1997] S. A. Empedocles and M. G. Bawendi, Quantum-confined stark effect in single cdse nanocrystallite quantum dots, Science 278, 2114 (1997).
- Xu et al. [2008] X. Xu, A. Andreev, and D. A. Williams, Manipulating quantum-confined stark shift in electroluminescence from quantum dots with side gates, New Journal of Physics 10, 053036 (2008).
- Bozyigit et al. [2013] D. Bozyigit, O. Yarema, and V. Wood, Origins of low quantum efficiencies in quantum dot leds, Advanced Functional Materials 23, 3024 (2013).
- Salihoglu et al. [2016] O. Salihoglu, N. Kakenov, O. Balci, S. Balci, and C. Kocabas, Graphene as a reversible and spectrally selective fluorescence quencher, Scientific Reports 6, 33911 (2016).
- Chen et al. [2012] R. Chen, Y. Gao, G. Zhang, R. Wu, L. Xiao, and S. Jia, Electric field induced fluorescence modulation of single molecules in pmma based on electron transfer, International Journal of Molecular Sciences 13, 11130 (2012).
- Wang et al. [2013] S. Wang, E. B. Berda, X. Lu, X. Li, C. Wang, and D. Chao, Tuning the fluorescent response of a novel electroactive polymer with multiple stimuli, Macromolecular Rapid Communications 34, 1648 (2013).
- Xie et al. [2022] S. Xie, H. Zhu, M. Li, and V. Bulović, Voltage-controlled reversible modulation of colloidal quantum dot thin film photoluminescence, Applied Physics Letters 120 (2022).
- Leng et al. [2018] H. Leng, B. Szychowski, M.-C. Daniel, and M. Pelton, Strong coupling and induced transparency at room temperature with single quantum dots and gap plasmons, Nature Communications 9, 4012 (2018).
- Limonov et al. [2017] M. F. Limonov, M. V. Rybin, A. N. Poddubny, and Y. S. Kivshar, Fano resonances in photonics, Nature Photonics 11, 543 (2017).
- Tasgin et al. [2018] M. E. Tasgin, A. Bek, and S. Postacı, Fano resonances in the linear and nonlinear plasmonic response, Fano Resonances in Optics and Microwaves: Physics and Applications 219 (2018).
- Wu et al. [2010] X. Wu, S. K. Gray, and M. Pelton, Quantum-dot-induced transparency in a nanoscale plasmonic resonator, Optics Express 18, 23633 (2010).
- Fleischhauer et al. [2005] M. Fleischhauer, A. Imamoglu, and J. P. Marangos, Electromagnetically induced transparency: Optics in coherent media, Reviews of Modern Physics 77, 633 (2005).
- Shah et al. [2013] R. A. Shah, N. F. Scherer, M. Pelton, and S. K. Gray, Ultrafast reversal of a fano resonance in a plasmon-exciton system, Physical Review B—Condensed Matter and Materials Physics 88, 075411 (2013).
- Garrido Alzar et al. [2002] C. L. Garrido Alzar, M. A. G. Martinez, and P. Nussenzveig, Classical analog of electromagnetically induced transparency, American Journal of Physics 70, 37 (2002).
- Larocque et al. [2024] H. Larocque, M. A. Buyukkaya, C. Errando-Herranz, C. Papon, S. Harper, M. Tao, J. Carolan, C.-M. Lee, C. J. Richardson, G. L. Leake, et al., Tunable quantum emitters on large-scale foundry silicon photonics, Nature Communications 15, 5781 (2024).
- Miller et al. [1985] D. A. Miller, D. Chemla, T. Damen, A. Gossard, W. Wiegmann, T. Wood, and C. Burrus, Electric field dependence of optical absorption near the band gap of quantum-well structures, Physical Review B 32, 1043 (1985).
- Rajput and Shishodia [2020] P. Rajput and M. S. Shishodia, Förster resonance energy transfer and molecular fluorescence near gain assisted refractory nitrides based plasmonic core-shell nanoparticle, Plasmonics 15, 2081 (2020).
- Issah et al. [2023] I. Issah, J. Pietila, T. Kujala, M. Koivurova, H. Caglayan, and M. Ornigotti, Epsilon-near-zero nanoparticles, Physical Review A 107, 023501 (2023).
- Murata et al. [2011] K. Murata, Y. Yasutake, K.-i. Nittoh, S. Fukatsu, and K. Miki, High-density g-centers, light-emitting point defects in silicon crystal, AIP Advances 1 (2011).
- Rotem et al. [2007] E. Rotem, J. M. Shainline, and J. M. Xu, Enhanced photoluminescence from nanopatterned carbon-rich silicon grown by solid-phase epitaxy, Applied Physics Letters 91 (2007).
- Ngandeu Ngambou et al. [2024] M. W. Ngandeu Ngambou, P. Perrin, I. Balasa, A. Tiranov, O. Brinza, F. Bénédic, J. Renaud, M. Reveillard, J. Silvent, P. Goldner, J. Achard, and A. Tallaire, Hot ion implantation to create dense NV center ensembles in diamond, Applied Physics Letters 124, 134002 (2024).
- [37] Ansys/Lumerical, Fluorescence enhancement, https://optics.ansys.com/hc/en-us/articles/360042161033-Fluorescence-enhancement.
- Yücel et al. [2020] O. Yücel, S. Ateş, and A. Bek, Single-photon nanoantenna with in situ fabrication of plasmonic Ag nanoparticle at an hBN defect center, arXiv preprint arXiv:2003.13824 (2020).
- Liu et al. [2017] Y.-C. Liu, B.-B. Li, and Y.-F. Xiao, Electromagnetically induced transparency in optical microcavities, Nanophotonics 6, 789 (2017).
- Johnson and Christy [1972] P. B. Johnson and R.-W. Christy, Optical constants of the noble metals, Physical Review B 6, 4370 (1972).
- Palik [1998] E. D. Palik, Handbook of optical constants of solids, Vol. 3 (Academic press, 1998).
- Shibata et al. [2013] K. Shibata, H. Yuan, Y. Iwasa, and K. Hirakawa, Large modulation of zero-dimensional electronic states in quantum dots by electric-double-layer gating, Nature Communications 4, 2664 (2013).
- Tasgin [2017] M. E. Tasgin, Many-particle entanglement criterion for superradiantlike states, Physical Review Letters 119, 033601 (2017).
- Alicki and Fannes [2013] R. Alicki and M. Fannes, Entanglement boost for extractable work from ensembles of quantum batteries, Physical Review E—Statistical, Nonlinear, and Soft Matter Physics 87, 042123 (2013).
- Ferraro et al. [2018] D. Ferraro, M. Campisi, G. M. Andolina, V. Pellegrini, and M. Polini, High-power collective charging of a solid-state quantum battery, Physical Review Letters 120, 117702 (2018).
- Bluvstein et al. [2024] D. Bluvstein, S. J. Evered, A. A. Geim, S. H. Li, H. Zhou, T. Manovitz, S. Ebadi, M. Cain, M. Kalinowski, D. Hangleiter, et al., Logical quantum processor based on reconfigurable atom arrays, Nature 626, 58 (2024).
- Lodahl et al. [2004] P. Lodahl, A. Floris van Driel, I. S. Nikolaev, A. Irman, K. Overgaag, D. Vanmaekelbergh, and W. L. Vos, Controlling the dynamics of spontaneous emission from quantum dots by photonic crystals, Nature 430, 654 (2004).
- Stürner et al. [2021] F. M. Stürner, A. Brenneis, T. Buck, J. Kassel, R. Rölver, T. Fuchs, A. Savitsky, D. Suter, J. Grimmel, S. Hengesbach, et al., Integrated and portable magnetometer based on nitrogen-vacancy ensembles in diamond, Advanced Quantum Technologies 4, 2000111 (2021).
- Madsen et al. [2022] L. S. Madsen, F. Laudenbach, M. F. Askarani, F. Rortais, T. Vincent, J. F. Bulmer, F. M. Miatto, L. Neuhaus, L. G. Helt, M. J. Collins, et al., Quantum computational advantage with a programmable photonic processor, Nature 606, 75 (2022).

### v2

@@
-Continuous and Reversible Electrical Tuning of Fluorescent Decay Rate via Fano Resonance
+Continuous and reversible electrical-tuning of fluorescent decay rate via Fano resonance
 Abstract
-Decay rate of an atomic or molecular dipole depends on the local density of optical states (LDOS) —Purcell effect—which can significantly be enhanced near a dielectric-core metal-shell nanoparticle (CSNP).
-On top of that, an auxiliary quantum object (QO) can introduce a Fano transparency in the plasmonic spectrum of the CSNP.
-Here, we show that an auxiliary QO, located at the hotspot of the CSNP, can modify the LDOS, hence the decay rate of an excited dipole.
-Moreover, by controlling the resonance of the auxiliary QO via an applied voltage, one can continuously and reversibly tune both the radiative and nonradiative decay rates of the dipole up to 2 orders-of-magnitude.
-This phenomenon emerges as an invaluable tool to implement in integrated quantum technologies, enabling realization of on-demand entanglement/single-photon sources, controlled execution of quantum gates and electrical-control of superradiant-like phase transitions.
-It also bears potential for application in super-resolution microscopy and surface-enhanced Raman spectroscopy (SERS).
+We demonstrate that the decay rates of a fluorescent molecule can be controlled by electrically shifting a transparency introduced by a Fano resonance.
+An auxiliary quantum object (QO), located at the hotspot of a plasmonic nanoparticle, suppresses plasmonic excitation at its level spacing .
+As a result, the local density of states (LDOS) associated with the plasmonic spectrum is also suppressed at .
+By shifting via an applied voltage, we continuously tune the radiative and nonradiative decay rates of the fluorescent molecule by up to two orders of magnitude.
+This mechanism offers a valuable tool for integrated quantum technologies, enabling on-demand entanglement and single-photon sources, voltage-controlled quantum gate operations, and electrical control of superradiant-like phase transitions.
+The approach also holds promise for applications in super-resolution microscopy and surface-enhanced Raman spectroscopy (SERS).
 I Introduction
@@
 Therefore, achieving continuous and reversible control within a compact device geometry that provides large modulation depths with (and over) CPU clock speed compatibility remains as a great challenge in active control of fluorescence for technology applications.
-In this paper, we provide a proof-of-principle demonstration of active and reversible tuning of both radiative and nonradiative decay rates of a fluorescent molecule (FM).
-We electrically tune the LDOS by shifting the position of a Fano transparency (also known as Fano resonance) [23, 24, 25].
-The resonance of the Fano transparency, induced by a quantum object (QO) residing at the CSNP hotspot (Fig.
-1a), is tuned by shifting the level-spacing of the QO via an applied voltage.
-This way, the decay rates can be continuously tuned between =1–215 within a “picosecond” response time which is limited by the readout rate of the circuit operation (GB/s).
+While certain progresses have been achieved on the electrical control of the Purcell effect [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22], they either remain below the CPU clock speeds or possess poor modulation depths.
+In this paper, we introduce a new method for fast electrical tuning of the Purcell effect that also achieves large modulation depths.
+Unlike conventional methods used in the literature [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22], our approach relies on tuning a transparency in the plasmonic response that is introduced via a Fano resonance [23, 24, 25].
+Our method outperforms existing electric-control techniques by offering a modulation depth of approximately 200 (i.e., 20,000%) and a response time on the order of a picosecond.
+It allows continuous tuning of the radiative and nonradiative decay rates of a fluorescent molecule, or the control of spontaneous parametric down-conversion in quantum integrated circuits.
+An auxiliary quantum object (QO) —such as a quantum dot or a collection of defect centers [26, 27, 28]— is placed at the hotspot of a core-shell nanoparticle (CSNP) and introduces the Fano resonance (i.e., a transparency) [23, 24, 25], as shown in Fig.
+1.
+This transparency appears at the level-spacing of the QO, which is = 766 nm in our simulations 111We refer the frequencies by the corresponding free-space wavelength, governed by .
+.
+The transparency suppresses plasmonic excitation and, in turn, eliminates the plasmonic enhancement in the local density of states (LDOS) at (see Fig.
+2).
+When the metastable transition frequency of a fluorescent molecule —located 15 nm away— matches the QO level spacing (i.e., ), the plasmonic enhancement of fluorescence is reset to its vacuum value, (compare solid and dashed lines in Fig.
+2).
+The QO level spacing—and hence the spectral position of the transparency—can be controlled by an applied voltage (see Fig.
+3a).
+For example, if shifts to = 767 nm, the new transparency appears at 767 nm.
+As a result, the LDOS associated with the decay at nm changes from to .
+By continuously shifting the QO level spacing, the decay rate can be tuned smoothly (see Fig.
+3b and 3c).
+This method allows full control over the range to by adjusting the voltage from to 1 V.
+This mechanism enables our approach to achieve both large modulation depths and fast response times, limited only by the circuit readout rate (GB/s).
 We emphasize that the demonstrated system is fundamentally different from previously reported active PL switching and modulation effects that are based on screening of the PL for an external observer.
 By trapping the fluorescent state in the molecule (i.e., not quenching it), the nature of our switching process exhibits a crucial property with the ability to hold and release the interested fluorescent state by voltage application.
+We provide a proof-of-principle demonstration of active and reversible tuning of both radiative and nonradiative decay rates of a fluorescent molecule (FM) via exact solutions of the 3D Maxwell equations.
+We perform our simulations using in Lumerical –a reliable FDTD simulation toolbox [30].
 In our demonstrated system, the presence of a plasmonic SiO2/Ag core (SiO2) shell (Ag) nanoparticle (CSNP) boosts the LDOS of a FM (Fig.
 1a).
-When an auxiliary QO, such as a quantum dot (QD) or a defect-center of resonance is positioned near the CSNP, it introduces a Fano transparency at [26].
+When an auxiliary QO, such as a quantum dot (QD) or a defect-center of resonance , is positioned near the CSNP, it introduces a Fano transparency at [31].
 Fano transparency is the plasmon-analog of electromagnetically-induced-transparency (EIT) [32].
@@
 Therefore, an absorption and scattering dip appears at .
+In Sec.
+I of the Supplementary Material (SM) [35], we explain why and how such a transparency arises in the plasmonic excitation spectrum, using a single equation with detailed discussion.
 In this work, by choosing about the transition frequency of the FM’s metastable (fluorescence) state (Fig.
@@
 This, however, won’t change the nature of the demonstrated phenomenon.
+Moreover, in the supplementary material (SM) we demonstrate that such a control mechanism works for different choices of the FM-MNP distances (see Sec.
+IV of the Supplementary Material (SM) [35]).
 III Conclusion, Discussion and Outlook
-We provide a proof-of-principle numerical demonstration for the voltage-tunability of decay rates.
-Our simulations —exact solutions of Maxwell equations— show 2-orders magnitude modulation in the decay rate.
-A QO turns off the CSNP induced LDOS at fluorescence = .
-When is shifted by an applied voltage (Fig.
-3a), the schemes (Figs.
-2a and 2b) for the decay rates also shift, enabling a continuous tuning of these rates in Figs.
-3b and 3c.
+We present a proof-of-principle numerical demonstration of voltage-tunable decay rates.
+Unlike conventional methods [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22], we utilize a transparency introduced by a Fano resonance.
+This transparency occurs at the level spacing of an auxiliary quantum object (QO) placed at the hotspot of a plasmonic nanoparticle (see Fig.
+1).
+We control the decay rate of a fluorescent molecule located 15 nm away from the nanoparticle.
+The transparency suppresses the LDOS induced by the nanoparticle (CSNP), which in turn reduces the decay rate of the fluorescent molecule.
+By shifting via an applied voltage, we control the position of the transparency and, consequently, the decay rate continuously (see Figs.
+3b and 3c).
+Our simulations—based on exact solutions of the 3D Maxwell equations—demonstrate up to two orders of magnitude modulation in the decay rate.
 The phenomenon addresses a wide range of applications as we are demonstrating a fundamental issue in light-matter interaction, that is electrical-tuning of LDOS.
 This provides an invaluable tool especially for integrated quantum technologies.The technique offers ultimate control over on-demand single-photon sources and entanglement generation in circuits.
+The method can also be utilized to control spontaneous parametric down conversion rate by choosing at .
 These also imply voltage-controlled quantum gate operations.
@@
 We anticipate that the demonstrated technique can also be utilized in quantum optics for the on-demand generation of single-photons (or photon-pairs) and their synchronization with external systems such as a vibrating scanning probe microscopy (SPM) tip, or a confocal scanning laser microscope.
+Unlike electrically-triggered quantum dots or defects [50], our method tunes the mean decay time after electrical excitation, and it can also be combined with such sources.
 Thanks to the ultimate control on the time gated activation of the emitter, the technique can also prove to be useful in performing spin-echo based quantum measurements/sensing applications such as quantum magnetometers [51] or quantum information processing [52].
@@
 Postacı, Fano resonances in the linear and nonlinear plasmonic response, Fano Resonances in Optics and Microwaves: Physics and Applications 219 (2018).
+- Murata et al. [2011] K. Murata, Y. Yasutake, K.-i. Nittoh, S. Fukatsu, and K. Miki, High-density g-centers, light-emitting point defects in silicon crystal, AIP Advances 1 (2011).
+- Rotem et al. [2007] E. Rotem, J. M. Shainline, and J. M. Xu, Enhanced photoluminescence from nanopatterned carbon-rich silicon grown by solid-phase epitaxy, Applied Physics Letters 91 (2007).
+- Ngandeu Ngambou et al.
+[2024] M.
+W.
+Ngandeu Ngambou, P.
+Perrin, I.
+Balasa, A.
+Tiranov, O.
+Brinza, F.
+Bénédic, J.
+Renaud, M.
+Reveillard, J.
+Silvent, P.
+Goldner, J.
+Achard, and A.
+Tallaire, Hot ion implantation to create dense NV center ensembles in diamond, Applied Physics Letters 124, 134002 (2024).
+- Note [1] We refer the frequencies by the corresponding free-space wavelength, governed by .
+- [30] Ansys/Lumerical, Fluorescence enhancement, https://optics.ansys.com/hc/en-us/articles/360042161033-Fluorescence-enhancement.
 - Wu et al. [2010] X. Wu, S. K. Gray, and M. Pelton, Quantum-dot-induced transparency in a nanoscale plasmonic resonator, Optics Express 18, 23633 (2010).
@@
 - Garrido Alzar et al. [2002] C. L. Garrido Alzar, M. A. G. Martinez, and P. Nussenzveig, Classical analog of electromagnetically induced transparency, American Journal of Physics 70, 37 (2002).
+- [35] See Supplemental Material at ”https://”.
 - Larocque et al.
@@
 - Issah et al. [2023] I. Issah, J. Pietila, T. Kujala, M. Koivurova, H. Caglayan, and M. Ornigotti, Epsilon-near-zero nanoparticles, Physical Review A 107, 023501 (2023).
-- Murata et al. [2011] K. Murata, Y. Yasutake, K.-i. Nittoh, S. Fukatsu, and K. Miki, High-density g-centers, light-emitting point defects in silicon crystal, AIP Advances 1 (2011).
-- Rotem et al. [2007] E. Rotem, J. M. Shainline, and J. M. Xu, Enhanced photoluminescence from nanopatterned carbon-rich silicon grown by solid-phase epitaxy, Applied Physics Letters 91 (2007).
-- Ngandeu Ngambou et al.
-[2024] M.
-W.
-Ngandeu Ngambou, P.
-Perrin, I.
-Balasa, A.
-Tiranov, O.
-Brinza, F.
-Bénédic, J.
-Renaud, M.
-Reveillard, J.
-Silvent, P.
-Goldner, J.
-Achard, and A.
-Tallaire, Hot ion implantation to create dense NV center ensembles in diamond, Applied Physics Letters 124, 134002 (2024).
-- [37] Ansys/Lumerical, Fluorescence enhancement, https://optics.ansys.com/hc/en-us/articles/360042161033-Fluorescence-enhancement.
 - Yücel et al. [2020] O. Yücel, S. Ateş, and A. Bek, Single-photon nanoantenna with in situ fabrication of plasmonic Ag nanoparticle at an hBN defect center, arXiv preprint arXiv:2003.13824 (2020).
@@
 Vos, Controlling the dynamics of spontaneous emission from quantum dots by photonic crystals, Nature 430, 654 (2004).
+- Yuan et al.
+[2002] Z.
+Yuan, B.
+E.
+Kardynal, R.
+M.
+Stevenson, A.
+J.
+Shields, C.
+J.
+Lobo, K.
+Cooper, N.
+S.
+Beattie, D.
+A.
+Ritchie, and M.
+Pepper, Electrically driven single-photon source, science 295, 102 (2002).
 - Stürner et al.